Please can you help me with this problem?
sqrt 30x-15 = x+7

Answer by stanbon(75887)

sqrt(30x-15) = x+7
Square both sides to get:
30x-15 = x^2+14x+49
x^2-16x+64=0
(x-8)^2 = 0
x = 8
===============
Cheers,
Stan H.